### v2

# Causal Discovery in Financial Markets: A Framework for Nonstationary Time-Series Data

Abstract
A deeper comprehension of financial markets necessitates understanding not only the statistical dependencies among various entities but also the causal dependencies. This paper extends the Constraint-based Causal Discovery from Heterogeneous Data algorithm to account for lagged relationships in time-series data (an algorithm we call CD-NOTS), shedding light on the complex causal relations between different financial assets and variables. We compare the performance of different algorithmic choices, such as the choice of conditional independence test, to give general advice on the effective way to use CD-NOTS. Using the results from the simulated data, we apply CD-NOTS to a broad range of indices and factors in order to identify causal connections among the entities, thereby showing how causal discovery can serve as a valuable tool for factor-based investing, portfolio diversification, and comprehension of market dynamics. Further, we show our algorithm is a more effective alternative to other causal discovery algorithms since the assumptions of our algorithm are more realistic in terms of financial data, a conclusion we find is statistically significant.
1 Introduction
Actions have consequences; that explains causality in simple words. Going beyond correlation or association, it is a guide which helps us gain a deeper understanding of the relationships between events, allowing us to perceive the impact of one phenomenon on another. An empirical causal framework would consist of three phases: (a) causal discovery; (b) causal inference; (c) causal explainability. In the first phase, the causal network is estimated from the data based on independence tests. When the network is discovered, then different scenarios can be tested on the structure. Afterwards, we can dive deep into the intuitive why of the causal network and results.
In this paper, our focus is on causal discovery, which aims to uncover the causal dependency structure from the observed data (Spirtes et al., 2001). While the traits in the data narrow down the possible causal network structures, it may not necessarily be sufficient to discover the complete underlying true causal network.
Most causal discovery methods can be categorized into three types of algorithms: constraint-based (e.g., Spirtes et al. (2000); Huang et al. (2020)), score-based (e.g., Chickering (2003); Silander and Myllymäki (2012); Lam et al. (2022)), and functional causal model-based (FCM) (e.g., Shimizu et al. (2006, 2011); Hyvärinen et al. (2010)). The constraint-based causal discovery algorithm infers causal dependency between nodes by running statistical independence tests among conditional and marginal probabilities to establish the network structure. The score-based algorithm defines a score criterion (e.g., likelihood scores Schwarz (1978); Buntine (1991)) and optimizes for the causal structure that best fits the data. On the other hand, FCMs intend to model the functional mechanisms of the variables. This calls for additional assumptions on the distribution of the data or domain-specific knowledge.
Causal discovery algorithms require assumptions about the data generating process; thus, each algorithm has limitations depending on the realism of the assumptions. For example, some only work on linear systems (Granger, 1969), some assume a Gaussian distribution for the variables (Ramsey and Andrews, 2017), and many assume there are no latent (unobserved) confounders (Zhang and Hyvärinen, 2009). Further, implementing causal discovery algorithms on time-series data adds additional challenges, such as sampling interval and nonstationarity, to the existing issues. There are ways to reduce the nonstationarity of time-series data such as using stock returns as opposed to price. However, empirically, we find this is insufficient to make the process stationary. Unlike evaluation of predictive performance, in general, causal discovery algorithms, when applied in practice, do not have ground truth data or ways to test the accuracy of the discovered network via randomized controlled trials. This leads to the importance of ensuring that the assumptions of the causal discovery algorithm pertains to the situations in which they are used. In our case studies (Section 5), we show that, in the financial examples we experiment with, the data is neither linear, Gaussian, nor stationary.
Refer to captionptCD-NOTSRefer to caption
Figure 1: We show a high-level diagram of the problem we are solving. We insert some time-series data into our causal discovery algorithm, from which we get a causal diagram of how the variables are related.
Our developed algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), overcomes the limitations mentioned above. It is a nonparametric approach, meaning it can capture both linear and non-linear causal relations and can also capture non-Gaussian distributions. It is capable of handling nonstationarity, and it is able to detect both lagged and contemporaneous causalities. Prior work has developed and used causal discovery algorithms that capture subsets of these properties; our work is the first (to the best of our knowledge) to tackle all three. We then analyze the patterns found in financial data using this algorithm.
The structure of this paper is as follows. In Section 2, we outline the related work and literature. Moving on to Section 3, we explain our methodology in detail. In Section 4, we conduct an investigation of simulated and real financial data to assess the framework’s validity and performance. Finally, in Section 5, we summarize the key findings of the paper and propose potential avenues for future research.
The most similar to our work is Huang et al. (2020), which we extended to handle lagged relationships in time-series data. In Section 3.3, we show a simple example in which we illustrate the difference in the causal relationships discovered when accounting for lagged relationships. Huang et al. (2020) discusses in their work a way in which their algorithm can be extended to lagged relationships; however, in Section 3.4, we show a simple example in which their extension would discover an incorrect graph.
2 Related Work
Causality explores the mechanisms that influence the outcomes, providing a more comprehensive picture of the why and how of unfolding events. The Rubin Causal Model (Rubin, 1974) and Pearl’s Structural Causal Model (SCM) (Pearl, 2009) are two approaches that are mainly considered when talking about causal relations.. A Structural Causal Model (Pearl, 2014) defines the causal relationships between variables via functions, taking the cause and the effect variables as input and output respectively. The SCM can be represented as a graphical model which can then be used for causal inference. On the other hand, the Rubin model (Splawa-Neyman et al., 1990), also known as the Potential Outcomes Framework, defines the conditional independence properties of the treatment assignment mechanism in order to analyze the unit-based response variable. Often, this framework requires a set of covariates that satisfy conditional ignorability, or in other words, no unobserved confounders. Note that prior work has shown that SCMs are formally equivalent to the Potential Outcomes Framework (Pearl, 2009).
Many statisticians have formulated multiple computational techniques and procedures to discover causal links through learning the causal relationships, typically depicted as directed acyclic graphs, as reviewed in Glymour et al. (2019). After obtaining the causal skeleton, there is an attempt to orient the edges. The two main methods are constraint-based and score-based. Constraint-based ones can handle confounders as the score-based approach is applicable only when there is a no-confounder assumption.
Constraint-based algorithms run independence-based tests to establish the underlying causal network (Nogueira et al., 2022). This set of algorithms starts from a full graph, orients edges based on conditional independence (CI), and identifies colliders. It is an iterative process which ends when there are no further edges to orient. Two widely used algorithms are Peters and Clark (PC), and Fast Causal Inference (FCI) (Spirtes et al., 2000). PC assumes the absence of confounders, which results in asymptotically correct causal information that helps with simplifying the problem. There have been improvements to the PC algorithm that tend to reduce the number of conditional independence tests and relax the assumptions (e.g., Tsamardinos et al. (2006); Bühlmann et al. (2009); Ramsey et al. (2012)). On the other hand, FCI produces asymptotically correct outcomes even when confounders are present; Colombo et al. (2012) and Rohekar et al. (2021) are two improved FCI works. Since our work is an extension of the PC algorithm, we utilize four different tests: Kernel-based Conditional Independence Test (KCIT) (Zhang et al., 2012), Randomized Conditional Correlation Test (RCoT) (Strobl et al., 2017), Conditional Mutual Information K-nearest neighbor (CMIknn) Test (Runge, 2017), and Partial Correlation (ParCorr) (Baba et al., 2004).
More recently, causal structure discovery algorithms have been developed that do not require statistical tests but instead use assumptions of the functional form or of the distribution in order to recover the structure (Shimizu et al., 2006; Zhang and Hyvärinen, 2009, 2010). For example, these algorithms assume absence of confounders and a non-Gaussian distribution within the data, enabling them to uncover the asymmetry inherent in the causes and effects distribution (Gendron et al., 2023).
Causal discovery in time-series data is about uncovering and quantifying causal relationships among variables that evolve over time. The difference between causal discovery methods and traditional time-series analysis is that causal discovery aims to identify the cause and effect relationships in the data, while traditional time-series analysis is more focused on prediction.
Granger causality (Granger, 1969) is an approach that estimates the linear causal relations using a VAR model (Hyvärinen et al., 2010). There is additional work for causal structure recovery using Dynamic Bayesian Networks and Hidden Markov models (Moraffah et al., 2021). Another class of causal discovery methods uses information theoretic quantities such as Mutual Information (Palus et al., 2001), and Transfer Entropy (Schreiber, 2000). Furthermore, graphical approaches like Time-Series Fast Causal Inference (Entner and Hoyer, 2010) and PCMCI (Runge et al., 2019) are more recent graphical approaches that are tweaked to be able to handle time-series data.
Most of the time-series causal discovery literature mentioned are focused on stationary time-series, and their approaches might lead to spurious causal relations if the time-series contains distributional shifts. Some work has been developed to address this gap (e.g., Peters et al. (2015); Zhang et al. (2017)). One approach is to assume the consistency of the nonstationarity reflected in the network across time. For instance, Huang et al. (2020) proposes an algorithm for heterogeneous time-series that only checks the nonstationarity at one time $t$. Our developed algorithm extends this to handle nonstationarity in lagged relationships.
3 Methodology
Time-series is the fundamental type of data used in many domains, from finance to climate science and engineering, requiring customizing algorithms to take this type of data as input. This tailoring requires taking into account some of the characteristics that this type of data carries, such as temporality, lead-lag dependency, and nonstationarity. Similar to CD-NOD, our algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), has four steps: adding the time-indexed node, using conditional independence tests to discover the causal skeleton (undirected causal graph), orienting the edges via prior knowledge and specific structures (e.g., the arrow of time and V-structures), and, finally, orienting remaining edges based on causal change independence. By design, our framework is able to handle time-series as input and, accordingly, takes into account its unique characteristics; in comparison, CD-NOD does not handle lagged relationships (Section 3.4).
With suitable assumptions, the causal network, using only observational data, can be estimated using conditional independence tests (Schölkopf et al., 2021). We follow the same set of assumptions commonly used in the causal network field.
Assumption 1 (Causal Sufficiency).
We assume that any potential confounders can be expressed as a smooth function of time. Hence, the time index node is a representation of all confounders.
Assumption 2 (Causal Faithfulness).
We assume that the conditional independence between the observed data is a true reflection of the absence of a direct relation in the variables given the conditional set. In other words, if $X$ is conditionally independent of $Y$ given $Z$ ($X\perp\!\!\!\perp Y|Z$), then this is true in the underlying causal network.
Assumption 3 (Randomness).
The data points are randomly selected from the population implied by the causal model (i.e., there is no selection bias).
Assumption 4 (Causal Consistency).
We assume that the causal relations between the variables, including the time indexed node, are consistent through time. This means that when lagged relations are present, the contemporaneous causal relations are the same in each lag and the cross lag causal relations also repeat themselves.
3.1 CD-NOTS Algorithm
We extend the algorithm described in Huang et al. (2020) in order to identify the causal network for time-series. Our developed approach, termed Constraint-based Causal Discovery from Nonstationary Time-Series (CD-NOTS), enhances the original method for identifying causal relationships between time-series. We show a high-level diagram of our algorithm in Figure 2.
Stage 1
Assume we have $N$ time-series. Within our network setup, each time-series is treated as a distinct node, denoted as $V_{i,t}$, representing the time-series $i$ at time $t$. To handle nonstationarity, we introduce a time-indexed node, $U_{t}$. Specifically, for any nonstationary time-series $i$, we anticipate a connection between $V_{i,t}$ and $U_{t}$ across all time points. We start from a full graph including the time-indexed node, where all the edges are drawn out.
Stage 2
With Assumptions 1, 2 and 3 in place, we reconstruct the causal skeleton. Employing CI tests, we investigate the independence between a node and the time-indexed node given a subset of nodes $S$. This process figures out the edges between nodes $U_{t}$ and $V_{i,t}$. The hypothesis test is:
$H_{0}\mathrel{\mathop{:}}P(V_{i,t},U_{t}|S)=P(V_{i,t}|S)\,P(U_{t}|S)\,,$
$H_{1}\mathrel{\mathop{:}}P(V_{i,t},U_{t}|S)\neq P(V_{i,t}|S)\,P(U_{t}|S)\,,$
which we test for all subsets $S\subseteq\{V_{k,t-l}\ |\ k=1,...,N;l=0,...,L\}\setminus\{V_{i,t}\}$ where $L$ denotes the maximum considered lag. Note that this hypothesis examines the relationship between the time-indexed node and other nodes at the identical time point $t$, maintaining consistency across time $t-l$.
Afterwards, utilizing the same test, we evaluate the independence of node pairs $V_{i,t-l}$ and $V_{j,t}$ excluding $U_{t}$. The null hypothesis is that the pair are conditionally independent, implying that their conditional probability equals the product of their marginals. If the obtained p-value falls below the confidence level, signifying dependence, the edge between the pair remains. The hypothesis test is:
$H_{0}\mathrel{\mathop{:}}P(V_{i,t-l},V_{j,t}|S)=P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
$H_{1}\mathrel{\mathop{:}}P(V_{i,t-l},V_{j,t}|S)\neq P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
where $S\subseteq\{U_{t}\}\cup\{V_{k,t-m}\ |\ k=1,...,N;m=0,...,L\}\setminus\{V_{i,t-l},V_{j,t}\}$. It is worth noting that when $l=0$, indicating no lag, we capture the contemporaneous relationships between nodes. On the other hand, when $l>0$, we discover the lagged dependencies among the time-series. Under Assumption 4, we maintain the structure consistent over time. This means that the discovered dependencies between $V_{i,t-l},V_{j,t}$ remain the same for $V_{i,t-l-m},V_{j,t-m};m=1,...,l-1$. The discovered network is undirected and has the edge minimality condition. This condition signifies that the network has the smallest number of causal changing mechanisms. This is due to the faithfulness across the entire network (Ghassami et al., 2018).
Stage 3
Once we have derived the network skeleton, our next step involves orienting the edges based on existing knowledge. If there are any nonstationary nodes, meaning there exists an edge between the node and time indexed node, then the orientation is from the time indexed node to the nonstationary node, as time is what causes the node’s corresponding time-series to be nonstationary. Moreover, all the edges where $l<m$ between pairs $V_{i,t-l}$ and $V_{j,t-m}$, the orientation is from $V_{i,t-m}$ to $V_{j,t-l}$, as logically the effect would happen after the cause. These directed edges along with the conditional independence tests can be used to orient other edges using the unique properties of V-structures (colliders) (Spirtes et al., 2000; Meek, 1995).
Stage 4
There might still remain some edges that are not yet directed. In the fourth stage, we rely on the independent changes of the causal modules (Pearl, 2009). A dependency measure is defined as an extension of the Hilbert Schmidt Independence Criterion (Gretton et al., 2008), which shows the level of dependency between causal modules. This is done by developing a kernel embedding of nonstationary conditional distributions and using their Gram matrix to create a test statistic (Huang et al., 2017). We apply the Meek orientation rule (Meek, 1995) on top of this to orient other edges. Note that there may remain edges that are undirected even after the third and fourth stage; in this scenario, no comments can be given on the causal direction of the corresponding pair of nodes.
[…]
Figure 2: CD-NOTS algorithm schematic example
3.2 Testing the Assumptions
The main assumptions we claim that are incorrectly assumed in prior causal discovery algorithms is stationarity, no lagged relationships, and linearity. Given the fact our algorithm does not make the first two assumptions, we can analyze the graph from our algorithm to see if nonstationary and lagged relationships are found; if found, then the conclusion is that, assuming stationarity (lagged relationships) is the null hypothesis, the statistical tests could not reject the null hypothesis.
On the other hand, for the linearity assumption, we modify an algorithm used in prior work (Peters et al., 2014) used in order to test post-hoc if the relationships are linear. Specifically, the relationships are assumed to be linear in conjunction assuming with additive noise. The algorithm used in Peters et al. (2014) is to perform a regression and then an independence test on the noise. Say the null hypothesis we are testing is that $X$ and $Y$ have a linear relationship given $Z$ (where we allow for $Z$ to have a non-linear relationship with $Y$ and the noise to be heteroscedastic with respect to $Z$), we perform a Gaussian process regression $\hat{f}$, where we use the linear kernel for $X$ and a non-linear kernel (e.g., RBF) for $Z$. Having done that, we then perform a CI test on $Y-\hat{f}(X,Y)\perp\!\!\!\perp X|Z$. Note that our test is more generic than testing if all relationships are linear.
3.3 Importance of Modeling Lagged Relationships
We show with a toy example how, if one were to not model lagged relationships but the data contains them, the graph would find nonstationary variables. Say we have one variable $X_{t}$ where:
$dX_{t}$ $=-x_{t}dt+\sqrt{2}dW_{t}\,,$
$X_{0}$ $=0\,.$
To discretize, we set $t$ to be integers. Since this is simply an Ornstein-Uhlenbeck (OU) process, we have:
$P(x_{t}|x_{u})$ $=P(x_{t}|\ \mathcal{N}(x_{u}e^{-\mathinner{\!\left\lvert t-u\right\rvert}},e^{-\mathinner{\!\left\lvert t-u\right\rvert}}))\,,\,u<t\,.$
From this, we can see that $P(x_{t}|x_{0})$ is a time dependent distribution whereas $P(x_{t}|x_{t-1})$ is not. This implies if we ran CD-NOTS assuming no lagged relationships (i.e., CD-NOD), $x_{t}$ would be considered nonstationary; whereas, if we ran CD-NOTS assuming a single lag, then $X_{t-1}\rightarrow X_{t}$ and $X_{t}$ would not be considered nonstationary.
3.4 Comparison to Algorithm in CD-NOD
[…]
Figure 3: Diagram of example causal graph for which CD-NOD’s time-series algorithm would make a mistake.
4 Evaluation
Figure 4: F-score evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
Figure 5: Runtime evaluation for CD-NOTS for different CI tests. Note the y-axis is in log scale.
In our algorithm, one major design decision is the conditional independence test. In order to ascertain which conditional independence test might be most effective, we test our algorithm on different simulated datasets. In total, we tested graphs with 3, 4, 5, 6, 8, 10, and 15 nodes, for a total of 350 graphs (50 graphs per number of nodes). For each graph, we tested having 50, 150, 300, 500, and 1,000 data points.
For the conditional independence tests, we specifically test using partial correlations (ParCorr, a conditional independence test that assumes linearity), Kernel-based Conditional Independence Test (KCIT), Randomized Conditional Correlation Test (RCoT), and Conditional Mutual Information K-nearest neighbor (CMIknn) Test. For both KCIT and RCoT, there are a few variations in how the p-value can be computed, specifically using the Satterthwaite–Welch (SW) method (Welch, 1938; Satterthwaite, 1946) or Hall– Buckley–Eagleson (HBE) method (Hall, 1983; Buckley and Eagleson, 1988). The goal of these algorithms is to approximate the CDF of a weighted sum of squared normals, which is the limiting behavior of the test statistic for KCIT and RCoT. The hyperparameters used for each test can be found in Appendix A.
In Figure 4, we compare the average F-score for each CI test across a varying number of nodes in the graph and a varying number of data points per graph. As expected, the F-score improves as the number of data points increase. Possibly most interesting is that ParCorr performs the best for low data regimes. Further, we can see, as found by previous work (Bodenham and Adams, 2015), that the SW variant of KCIT tends to be weaker than HBE (except when there are a large number of nodes and only 50 data points). However, there seems to be little difference for RCoT. So, relying on results in Bodenham and Adams (2015), we recommend the HBE variant. Finally, while CMIknn is quite competitive, we do not recommend its usage due to its lengthy runtime (Figure 5).
Comparing against the best tests (KCIT HBE, RCoT HBE, ParCorr, CMIknn), we consider PCMCI (Runge et al., 2019) as a benchmark since it is a comparable PC-based algorithm and is widely used in the causal discovery literature. The main difference in PCMCI over prior PC algorithms is its pruning of the set of variables which to condition on in the CI test. In Figure 6, we see that CD-NOTS consistently outperforms PCMCI.
Figure 6: F-score evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
5 Case Studies
We apply CD-NOTS on three different datasets: 1) Fama-French factors and Apple’s returns, 2) unemployment, CPI, and PPI of different countries, and 3) Price-to-Book ratio and stock returns of financial companies in the S&P 500. Most of the work done in factor investing and economic indices’ analysis rely on the correlation between the variables and do not consider the fundamental question of “why” (López de Prado, 2023). In each case study, we explain the data and properties we see in the data or has been found in prior work, discuss the implications of the causal networks found by CD-NOTS, and, finally, show that the assumptions of stationarity, linearity, and no lagged relationships would be incorrect to assume and thus, any algorithm making said assumptions would be inaccurate and invalid to use on these datasets.
For all the graphs in the section, we color by the maximum p-value for that edge across all tests run during the causal skeleton step (Stage 2). While we recognize that p-values are not a metric to measure the strength of the relationship, we can think of the vicinity of the p-value to 0.05 as a measure of the sensitivity of the graph to that hyperparameter (i.e., the threshold).
5.1 Fama-French Factors
Figure 7: Time-series of Fama-French factors and Apple’s returns.
Refer to caption Refer to caption Refer to caption Refer to caption (a) Jan 2000 - Dec 2007 Refer to caption (b) Jan 2008 - Dec 2015 Refer to caption (c) Jan 2016 - Dec 2022
Figure 8: Results of CD-NOTS KCIT run on three non-overlapping time periods.
Refer to caption Refer to caption
Figure 9: Results of CD-NOTS KCIT run from beginning of 2000 to end of 2022.
Data
For our first case study, we analyzed the relationship between the Fama-French factors (Fama and French, 1993) and Apple’s returns from the beginning of 2000 to the end of 2022. In the data in Figure 7, we show that there is some correlation between the variables (e.g., the spike in volatility around 2008 can be seen in all six subplots) and, accordingly, all variables show volatility clustering.
Results and Discussion
In Figure 8, we ran CD-NOTS using KCIT (RCoT results are in Appendix B) on three non-overlapping periods of six to seven years (approximately 2000 observations each). We note in all three periods that there is no connection from $T$ to Apple’s returns (AAPL_RET), implying that the nonstationarity observed in Apple’s returns can be explained away by the Fama-French factors. Further, the lagged relationships tends to be sparse, with many of the found lagged relationships having a high p-value (close to 0.05). However, we see that the relationship between the factors seems to change over time, as well as the relationship between the factors and Apple’s returns. In Figure 9, we run CD-NOTS with KCIT on the whole time range, and we can see now that that there is a connection from $T$ to Apple’s returns (AAPL_RET). This is to be expected for two reasons: 1) over the course of the last 22 years, Apple’s business model has changed, leading to exposures to different factors over time, and 2) the changing relationship between the factors and Apple’s returns, as seen in Figure 8.
Evaluating Assumptions
One of the assumptions of CD-NOTS is causal consistency that the causal graph does not change over time. We assert that the change in dependencies between the factors and Apple’s returns that we see in Figure 8 are not a result of changes in the causal directions but in the strength of the causal relationships as captured by the nonstationarity. Further, we note that the fact that there is a relationship between $T$ and all the Fama-French factors (nonstationarity) implies that any causal discovery algorithm assuming stationarity would give invalid results. Similarly, we find statistically significant lagged relationships. Further, for our linearity test, we test the linearity of Apple’s returns and each parent factor (CMA and Mkt-RF) and find the linearity hypothesis can be rejected for both factors; in other words, assuming linearity in the relationships between variables would be incorrect and thus could lead to incorrectly identified causal relationships.
5.2 Economic Data
Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
Figure 10: Results of CD-NOTS ParCorr run on six countries’ economic variables.
Figure 11: Results of CD-NOTS KCIT run on the economic variables of multiple countries.
Data
For our second case study, we analyzed the relationship between the month-over-month percentage change in CPI, the month-over-month percentage change in PPI, and unemployment for the U.S., Japan, Canada, India, Italy, the U.K., and France. We use monthly data from the beginning of 2000 to the end of 2023.
Results and Discussion
In Figure 10, we run CD-NOTS with ParCorr on the U.S., Japan, Canada, Italy, France, and the U.K.; we use ParCorr since we have no more than 300 observations per country. We can see that, while the graphs are not identical, they are quite similar, especially if we ignore edges with a higher p-value. Further, we notice the three graphs for France, the U.K., and Italy are very similar, possibly due to their geographic and economic proximity. We repeat the experiment with all countries using KCIT (Figure 11); further, we introduce another static variable (Country) into the graph. Similar to the previous case study in which the graph over all time periods is akin to a superset of the edges of individual time periods, we see a similar behavior here. Note that Country does not point to any other variable suggesting that the behavior across countries is independent of the country; this observation could be used to justify training a single econometric model across countries.
Evaluation Assumptions
Similar to the previous case study, we see statistically significant lagged relationships and statistically significant nonstationarity, implying we would not be able to trust the results of an algorithm that cannot handle these types of properties. For our linearity test, we test the linearity of the percentage change of CPI and each parent node (both including and excluding the contemporaneous percentage change in PPI since its direction was not inferred). We find that the only time the linearity hypothesis cannot be rejected is when using the contemporaneous percentage change in PPI as a parent and checking the linearity with the one-lag percentage change of CPI. We note that this is inconsistent with our usage of ParCorr in Figure 10 since ParCorr is a linear-based CI test. In Appendix C, we show the results of using KCIT and RCoT where we find qualitatively that the graphs are similar.
5.3 Company Financials and Returns
Refer to caption (a) Contemporaneous Effects Refer to caption (b) Lagged Effects Refer to caption (c) Contemporaneous Effects Refer to caption (d) Lagged Effects Refer to caption (e) Comparison of Returns and N. LPTB Refer to caption (f) Cumulative Returns of Strategies
Figure 12: Results of CD-NOTS KCIT run on financials data for log Price-to-Book ratio (LPTB) and stock returns (Ret). Normalized LPTB (N. LPTB) refers to normalizing the LPTB by the mean and standard deviation of the LPTB of financial companies in that quarter. For Figure 11(f), our strategy is long-ing financial companies that have a normalized LPTB greater than the threshold which we compare against the market-cap weighted portfolio.
Data
For our third case study, we analyzed the relationship between the Price-to-Book ratio and returns of financial companies in the S&P 500. We use data from the beginning of 2010 to the end of 2023. We utilize reported quarterly numbers, where we set the date to be the end of each quarter. To ensure no look-ahead bias when defining the Price-to-Book ratio, for any given date (i.e., end of quarter), we use the Price-to-Book ratio available that day. We join the next quarter returns with this company financials information. For our experiments, we use the log Price-to-Book ratio (LPTB) as well as the normalized LPTB, which we define as the LPTB normalized by the mean and standard deviation of the LPTB of financial companies in that quarter.
Results and Discussion
In Figure 11(b) and Figure 11(d), we compare the graphs we get running CD-NOTS using KCIT using normalized LPTB instead of LPTB. We find that the nonstationarity is removed (T does not point to N. LPTB). Further, we see that the previous quarter’s normalized LPTB point to the next quarter’s returns with a reasonable p-value suggesting that its inclusion in the graph is not purely a function of the threshold we chose. In Figure 11(e), we compare the next quarter returns and the normalized LPTB and see that the there is a small trend in the mean (the black circles). To this end, we create a simple trading strategy where we invest in the subset of the financial securities whose LPTB is greater than some threshold. In Figure 11(f), we see this simple approach outperforms a market-cap weighted strategy; prior work (Monge et al., 2023) has found that growth stocks have been outperforming value stocks. Importantly, we note that the goal of this is to not introduce a new signal but to show how causal discovery can be a tool in searching for statistically significant relations which, if the assumptions are true for the data, can be interpreted as causal.
Evaluating Assumptions
Similar to previous case studies, we see statistically significant lagged relationships and statistically significant nonstationarity, implying we would not be able to trust the results of an algorithm that cannot handle these types of properties. For our linearity test, we test the linearity of the normalized LPTB and returns conditional on all other parent nodes of returns and find the p-value be to 0.0 (i.e., we can reject the null hypothesis that the relationship is linear).
6 Conclusion and Future Work
In this paper, we developed a novel algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), which is nonparametric (meaning it can capture both linear and non-linear causal relations), can handle nonstationarity, and is able to detect both lagged and contemporaneous causalities. We showed using simulated data that, for effective use of CD-NOTS, one should use ParCorr for low data regimes (less than 200 data points) and KCIT or RCoT for higher data regimes.
We applied CD-NOTS to Fama-French factors and Apple’s returns, where we found that for small time ranges (approximately seven years), Apple’s returns are stationary conditioning on the factors. This justifies the usage of Fama-French factors as a method for returns attribution. However, over longer time ranges (approximately two decades), we find Apple’s returns are nonstationary even after accounting for the Fama-French factors, implying that there is a need to continually update the factor exposures of Apple through time.
Further, we found the causal relationship of economic factors (specifically unemployment, CPI, and PPI) tends to be similar across countries, especially for those geographically and economically similar. This finding implies that we should be able to create stronger models in economics by training a single model across different countries.
Finally, we apply CD-NOTS to company financials and stock returns and find a causal relationship between the Price-to-Book ratio and future returns of financial companies, justifying the common usage of Price-to-Book ratio for investing.
Besides our contribution in developing an algorithm that can handle nonstationarity, non-linear relationships, lagged and contemporaneous causalities, we showed the importance of these assumptions in finance. Specifically, many of the assumptions made in prior work (notably linearity, no lagged relationships, and stationarity) often do not hold in finance and thus should be used with caution. Hence, through weakening the assumptions required for our causal discovery algorithm to be used, we can trust more the causal relationships found by CD-NOTS and continue to explore the causal relationships in other financial datasets.
Appendix A Conditional Independence Tests Hyperparameters
For all conditional independence tests, we use a threshold of 0.05 for the p-value.
KCIT
For KCIT, we use the RBF kernel where the bandwidth is set to the median distance between the first five hundred data points.
RCoT
For RCoT, similar to KCIT, we use the RBF kernel where the bandwidth is set to the median distance between the first five hundred data points. Further, we use a fixed number of Fourier features, specifically 25 for $Z$ and 5 for $\ddot{X}$, $X$, and $Y$, similar to Strobl et al. (2017).
Appendix B Fama-French Factors Case Study with RCoT
We repeat the plots from our first case study (Section 5.1) using RCoT as opposed to KCIT. One empirical finding was that, due to the randomness of RCoT, many times there could be an inconsistency between the discovered graph and the CI tests, i.e., the graph shows conditional independences that are not found by the CI tests. We leave it to future work to find how to account for these forms on inconsistencies.
Refer to caption Refer to caption Refer to caption Refer to caption (a) Jan 2000 - Dec 2007 Refer to caption (b) Jan 2008 - Dec 2015 Refer to caption (c) Jan 2016 - Dec 2022
Figure 13: Results of CD-NOTS RCoT run on three non-overlapping time periods.
Refer to caption Refer to caption
Figure 14: Results of CD-NOTS RCoT run from beginning of 2000 to end of 2022.
Appendix C Economic Data Case Study with KCIT and RCoT
Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
Figure 15: Results of CD-NOTS KCIT run on six countries’ economic variables.
Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
Figure 16: Results of CD-NOTS RCoT run on six countries’ economic variables.
Figure 17: Results of CD-NOTS RCoT run on the economic variables of multiple countries.
Appendix D Detailed Simulation Results
D.1 Different CI Tests
Figure 18: F-score evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
Figure 19: Precision evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
Figure 20: Recall evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
Figure 21: Structural Hamming Distance evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
D.2 Comparison against PCMCI
Figure 22: F-score evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
Figure 23: Precision evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
Figure 24: Recall evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
Figure 25: Structural Hamming Distance evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.

### v3

@@
 Abstract
-A deeper comprehension of financial markets necessitates understanding not only the statistical dependencies among various entities but also the causal dependencies.
-This paper extends the Constraint-based Causal Discovery from Heterogeneous Data algorithm to account for lagged relationships in time-series data (an algorithm we call CD-NOTS), shedding light on the complex causal relations between different financial assets and variables.
-We compare the performance of different algorithmic choices, such as the choice of conditional independence test, to give general advice on the effective way to use CD-NOTS.
-Using the results from the simulated data, we apply CD-NOTS to a broad range of indices and factors in order to identify causal connections among the entities, thereby showing how causal discovery can serve as a valuable tool for factor-based investing, portfolio diversification, and comprehension of market dynamics.
-Further, we show our algorithm is a more effective alternative to other causal discovery algorithms since the assumptions of our algorithm are more realistic in terms of financial data, a conclusion we find is statistically significant.
+This paper introduces a new causal structure learning method for nonstationary time series data, a common data type found in fields such as finance, economics, healthcare, and environmental science.
+Our work builds upon the constraint-based causal discovery from nonstationary data algorithm (CD-NOD).
+We introduce a refined version (CD-NOTS) which is designed specifically to account for lagged dependencies in time series data.
+We compare the performance of different algorithmic choices, such as the type of conditional independence test and the significance level, to help select the best hyperparameters given various scenarios of sample size, problem dimensionality, and availability of computational resources.
+Using the results from the simulated data, we apply CD-NOTS to a broad range of real-world financial applications in order to identify causal connections among nonstationary time series data, thereby illustrating applications in factor-based investing, portfolio diversification, and comprehension of market dynamics.
+Keywords: Causal discovery, Structure learning, Nonstationary, Time series, Factor investing
 1 Introduction
 Actions have consequences; that explains causality in simple words.
-Going beyond correlation or association, it is a guide which helps us gain a deeper understanding of the relationships between events, allowing us to perceive the impact of one phenomenon on another.
+Going beyond correlation or association, it is a guide which helps us gain a deeper understanding of the relationships between events, allowing to perceive the impact of one phenomenon on another.
 An empirical causal framework would consist of three phases: (a) causal discovery; (b) causal inference; (c) causal explainability.
@@
 When the network is discovered, then different scenarios can be tested on the structure.
-Afterwards, we can dive deep into the intuitive why of the causal network and results.
-In this paper, our focus is on causal discovery, which aims to uncover the causal dependency structure from the observed data (Spirtes et al., 2001).
-While the traits in the data narrow down the possible causal network structures, it may not necessarily be sufficient to discover the complete underlying true causal network.
+Afterwards, we can dive deep into the why and explanations of the scenarios implied by the discovered network.
+In this paper, our focus is on causal discovery in time series, which aims to uncover the causal dependency structure from the observed data (Figure 1).
+Time series is the fundamental type of data used in many domains, from finance to climate science and engineering, requiring customized algorithms to take into account characteristics such as temporality, lead-lag dependency, and nonstationarity.
 Most causal discovery methods can be categorized into three types of algorithms: constraint-based (e.g., Spirtes et al.
@@
 (2010)).
-The constraint-based causal discovery algorithm infers causal dependency between nodes by running statistical independence tests among conditional and marginal probabilities to establish the network structure.
-The score-based algorithm defines a score criterion (e.g., likelihood scores Schwarz (1978); Buntine (1991)) and optimizes for the causal structure that best fits the data.
+The constraint-based causal discovery algorithm infers causal dependency between nodes by running statistical independence tests among conditional and marginal probabilities to establish the network structure (Tsagris et al., 2018; Nogueira et al., 2022).
+Two widely used algorithms of this type are Peters and Clark (PC), and Fast Causal Inference (FCI) (Spirtes et al., 2000).
+The score-based algorithm defines a score criterion (e.g., likelihood scores (Schwarz, 1978; Buntine, 1991)) and optimizes for the causal structure that best fits the data.
 On the other hand, FCMs intend to model the functional mechanisms of the variables.
@@
 Causal discovery algorithms require assumptions about the data generating process; thus, each algorithm has limitations depending on the realism of the assumptions.
-For example, some only work on linear systems (Granger, 1969), some assume a Gaussian distribution for the variables (Ramsey and Andrews, 2017), and many assume there are no latent (unobserved) confounders (Zhang and Hyvärinen, 2009).
-Further, implementing causal discovery algorithms on time-series data adds additional challenges, such as sampling interval and nonstationarity, to the existing issues.
-There are ways to reduce the nonstationarity of time-series data such as using stock returns as opposed to price.
+For example, some only work on linear systems (Granger, 1969), some assume a Gaussian distribution for the variables (Ramsey and Andrews, 2017), and many assume there are no latent (unobserved) confounders [^1] Some research has been conducted to first eliminate the impact of hidden variables and then apply the structural learning model (e.g., Frot et al.
+(2019))..(Zhang and Hyvärinen, 2009).
+Further, implementing causal discovery algorithms on time series data adds additional challenges, such as nonstationarity, to the existing issues.
+There are ways to reduce the nonstationarity of time series data such as using stock returns as opposed to prices.
 However, empirically, we find this is insufficient to make the process stationary.
 Unlike evaluation of predictive performance, in general, causal discovery algorithms, when applied in practice, do not have ground truth data or ways to test the accuracy of the discovered network via randomized controlled trials.
-This leads to the importance of ensuring that the assumptions of the causal discovery algorithm pertains to the situations in which they are used.
+This leads to the importance of ensuring that the assumptions of the causal discovery algorithm pertain to the situations in which they are used.
 In our case studies (Section 4), we show that, in the financial examples we experiment with, the data is neither linear, Gaussian, nor stationary.
-Refer to captionptCD-NOTSRefer to caption
-Figure 1: We show a high-level diagram of the problem we are solving.
-We insert some time-series data into our causal discovery algorithm, from which we get a causal diagram of how the variables are related.
-Our developed algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), overcomes the limitations mentioned above.
-It is a nonparametric approach, meaning it can capture both linear and non-linear causal relations and can also capture non-Gaussian distributions.
-It is capable of handling nonstationarity, and it is able to detect both lagged and contemporaneous causalities.
-Prior work has developed and used causal discovery algorithms that capture subsets of these properties; our work is the first (to the best of our knowledge) to tackle all three.
-We then analyze the patterns found in financial data using this algorithm.
-The structure of this paper is as follows.
-In Section 2, we outline the related work and literature.
-Moving on to Section 3, we explain our methodology in detail.
-In Section 4, we conduct an investigation of simulated and real financial data to assess the framework’s validity and performance.
-Finally, in Section 5, we summarize the key findings of the paper and propose potential avenues for future research.
-The most similar to our work is Huang et al.
-(2020), which we extended to handle lagged relationships in time-series data.
-In Section 3.3, we show a simple example in which we illustrate the difference in the causal relationships discovered when accounting for lagged relationships.
-Huang et al.
-(2020) discusses in their work a way in which their algorithm can be extended to lagged relationships; however, in Section 3.4, we show a simple example in which their extension would discover an incorrect graph.
-2 Related Work
-Causality explores the mechanisms that influence the outcomes, providing a more comprehensive picture of the why and how of unfolding events.
-The Rubin Causal Model (Rubin, 1974) and Pearl’s Structural Causal Model (SCM) (Pearl, 2009) are two approaches that are mainly considered when talking about causal relations..
-A Structural Causal Model (Pearl, 2014) defines the causal relationships between variables via functions, taking the cause and the effect variables as input and output respectively.
-The SCM can be represented as a graphical model which can then be used for causal inference.
-On the other hand, the Rubin model (Splawa-Neyman et al., 1990), also known as the Potential Outcomes Framework, defines the conditional independence properties of the treatment assignment mechanism in order to analyze the unit-based response variable.
-Often, this framework requires a set of covariates that satisfy conditional ignorability, or in other words, no unobserved confounders.
-Note that prior work has shown that SCMs are formally equivalent to the Potential Outcomes Framework (Pearl, 2009).
-Many statisticians have formulated multiple computational techniques and procedures to discover causal links through learning the causal relationships, typically depicted as directed acyclic graphs, as reviewed in Glymour et al.
-(2019).
-After obtaining the causal skeleton, there is an attempt to orient the edges.
-The two main methods are constraint-based and score-based.
-Constraint-based ones can handle confounders as the score-based approach is applicable only when there is a no-confounder assumption.
-Constraint-based algorithms run independence-based tests to establish the underlying causal network (Nogueira et al., 2022).
-This set of algorithms starts from a full graph, orients edges based on conditional independence (CI), and identifies colliders.
-It is an iterative process which ends when there are no further edges to orient.
-Two widely used algorithms are Peters and Clark (PC), and Fast Causal Inference (FCI) (Spirtes et al., 2000).
-PC assumes the absence of confounders, which results in asymptotically correct causal information that helps with simplifying the problem.
-There have been improvements to the PC algorithm that tend to reduce the number of conditional independence tests and relax the assumptions (e.g., Tsamardinos et al.
-(2006); Bühlmann et al.
-(2009); Ramsey et al.
-(2012)).
-On the other hand, FCI produces asymptotically correct outcomes even when confounders are present; Colombo et al.
-(2012) and Rohekar et al.
-(2021) are two improved FCI works.
-Since our work is an extension of the PC algorithm, we utilize four different tests: Kernel-based Conditional Independence Test (KCIT) (Zhang et al., 2012), Randomized Conditional Correlation Test (RCoT) (Strobl et al., 2017), Conditional Mutual Information K-nearest neighbor (CMIknn) Test (Runge, 2017), and Partial Correlation (ParCorr) (Baba et al., 2004).
-More recently, causal structure discovery algorithms have been developed that do not require statistical tests but instead use assumptions of the functional form or of the distribution in order to recover the structure (Shimizu et al., 2006; Zhang and Hyvärinen, 2009, 2010).
-For example, these algorithms assume absence of confounders and a non-Gaussian distribution within the data, enabling them to uncover the asymmetry inherent in the causes and effects distribution (Gendron et al., 2023).
-Causal discovery in time-series data is about uncovering and quantifying causal relationships among variables that evolve over time.
-The difference between causal discovery methods and traditional time-series analysis is that causal discovery aims to identify the cause and effect relationships in the data, while traditional time-series analysis is more focused on prediction.
-Granger causality (Granger, 1969) is an approach that estimates the linear causal relations using a VAR model (Hyvärinen et al., 2010).
-There is additional work for causal structure recovery using Dynamic Bayesian Networks and Hidden Markov models (Moraffah et al., 2021).
-Another class of causal discovery methods uses information theoretic quantities such as Mutual Information (Palus et al., 2001), and Transfer Entropy (Schreiber, 2000).
-Furthermore, graphical approaches like Time-Series Fast Causal Inference (Entner and Hoyer, 2010) and PCMCI (Runge et al., 2019) are more recent graphical approaches that are tweaked to be able to handle time-series data.
-Most of the time-series causal discovery literature mentioned are focused on stationary time-series, and their approaches might lead to spurious causal relations if the time-series contains distributional shifts.
+There is a line of approaches addressing the time series causal discovery problem such as Granger causality (Granger, 1969), Dynamic Bayesian Networks and Hidden Markov models (Moraffah et al., 2021), Mutual Information (Palus et al., 2001), Transfer Entropy (Schreiber, 2000), time series Fast Causal Inference (Entner and Hoyer, 2010), and PCMCI (Runge et al., 2019).
+Since most of the literature mentioned are focused on stationary time series and most current approaches do not handle heterogeneous variable types their approaches might lead to spurious causal relations if the time series contains distributional shifts.
 Some work has been developed to address this gap (e.g., Peters et al.
@@
 (2017)).
-One approach is to assume the consistency of the nonstationarity reflected in the network across time.
-For instance, Huang et al.
-(2020) proposes an algorithm for heterogeneous time-series that only checks the nonstationarity at one time $t$.
-Our developed algorithm extends this to handle nonstationarity in lagged relationships.
+Refer to captionCD-NOTSRefer to caption
+Figure 1: High-level diagram of the problem; we insert some time series data into our causal discovery algorithm, from which we get a causal diagram of how the variables are related.
+We design an algorithm for causal discovery from nonstationary time series, referred to as the “Constraint-based Causal Discovery from Nonstationary Time Series (CD-NOTS)” hereafter.
+This algorithm is an extension of the CD-NOD algorithm (Huang et al., 2020) to handle time series data [^2] Huang et al.
+(2020) discusses in their work a way in which their algorithm can be extended to lagged relationships; however, in Section 2.4, we show a simple example in which their extension would discover an incorrect graph..
+CD-NOTS handles nonstationarity as well as pseudo-confounders (defined in Section 2) while considering lagged relationships and enforces consistency of the graph through time.
+It is a nonparametric approach, meaning it can capture both linear and non-linear causal relations and also work with non-Gaussian distributions (e.g., Wang and Drton (2023)).
+In addition, it is able to detect both lagged and contemporaneous causal effects.
+The algorithm starts from a full graph, considering nodes for lagged time series, and designates a node to the time itself in order to model nonstationarity.
+Then, it uses Conditional Independence (CI) tests to uncover the causal skeleton.
+Leveraging background knowledge (e.g., Andrews et al.
+(2023)), discoverable structures, and causal change independence, it orients the edges.
+We experiment CD-NOTS with four different tests: Kernel-based Conditional Independence Test (KCIT) (Zhang et al., 2012), Randomized Conditional Correlation Test (RCoT) (Strobl et al., 2017), Conditional Mutual Information K-nearest neighbor (CMIknn) Test (Runge, 2017), and Partial Correlation (ParCorr) (Baba et al., 2004).
+The structure of this paper is as follows.
+In Section 2, we explain our methodology in detail.
+In Section 3, we conduct an investigation of simulated and real financial data to assess the framework’s validity and performance.
+Finally, in Section 4, we summarize the key findings of the paper and propose potential avenues for future research.
 2 Methodology
-Time-series is the fundamental type of data used in many domains, from finance to climate science and engineering, requiring customizing algorithms to take this type of data as input.
-This tailoring requires taking into account some of the characteristics that this type of data carries, such as temporality, lead-lag dependency, and nonstationarity.
-Similar to CD-NOD, our algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), has four steps: adding the time-indexed node, using conditional independence tests to discover the causal skeleton (undirected causal graph), orienting the edges via prior knowledge and specific structures (e.g., the arrow of time and V-structures), and, finally, orienting remaining edges based on causal change independence.
-By design, our framework is able to handle time-series as input and, accordingly, takes into account its unique characteristics; in comparison, CD-NOD does not handle lagged relationships (Section 3.4).
-With suitable assumptions, the causal network, using only observational data, can be estimated using conditional independence tests (Schölkopf et al., 2021).
+Following the CD-NOD approach, our algorithm, CD-NOTS, consists of four steps: adding the time-indexed node (which helps in handling nonstationarity), discovering the causal skeleton (undirected causal graph) using conditional independence tests, orienting the edges via prior knowledge (e.g., the arrow of time) and identifiable structures (e.g., V-structures) and, finally, orienting remaining edges based on causal change independence.
+With suitable assumptions, the causal network using only observational data, can be estimated by conditional independence tests (Schölkopf et al., 2021).
 We follow the same set of assumptions commonly used in the causal network field.
-Assumption 1 (Causal Sufficiency).
-We assume that any potential confounders can be expressed as a smooth function of time. Hence, the time index node is a representation of all confounders.
-Assumption 2 (Causal Faithfulness).
+Assumption 1 (Pseudo Causal Sufficiency)
+We assume that any potential confounders can be expressed as a smooth function of time. Hence, the time-indexed node is the representative of all the confounders.
+Assumption 2 (Causal Faithfulness)
 We assume that the conditional independence between the observed data is a true reflection of the absence of a direct relation in the variables given the conditional set.
 In other words, if $X$ is conditionally independent of $Y$ given $Z$ ($X\perp\!\!\!\perp Y|Z$), then this is true in the underlying causal network.
-Assumption 3 (Randomness).
+Assumption 3 (Randomness)
 The data points are randomly selected from the population implied by the causal model (i.e., there is no selection bias).
-Assumption 4 (Causal Consistency).
-We assume that the causal relations between the variables, including the time indexed node, are consistent through time.
-This means that when lagged relations are present, the contemporaneous causal relations are the same in each lag and the cross lag causal relations also repeat themselves.
+Assumption 4 (Causal Consistency)
+We assume that the causal relations between the variables, including the time-indexed node, are consistent through time.
+This means that when lagged relations are present, the contemporaneous causal relations are the same in each lag and the cross lag causal relations repeat themselves.
+We note that most prior work assumes sufficiency, i.e., cannot handle any unobserved confounders.
 2.1 CD-NOTS Algorithm
-We extend the algorithm described in Huang et al.
-(2020) in order to identify the causal network for time-series.
-Our developed approach, termed Constraint-based Causal Discovery from Nonstationary Time-Series (CD-NOTS), enhances the original method for identifying causal relationships between time-series.
-We show a high-level diagram of our algorithm in Figure 2.
+Our developed approach, CD-NOTS, identifies causal relationships between time series. We show a high-level diagram of our algorithm in Figure 2.
 Stage 1
-Assume we have $N$ time-series.
-Within our network setup, each time-series is treated as a distinct node, denoted as $V_{i,t}$, representing the time-series $i$ at time $t$.
+Assume we have $N$ time series.
+Within our network setup, each time series is treated as a distinct node, denoted as $V_{i,t}$, representing the time series $i$ at time $t$.
 To handle nonstationarity, we introduce a time-indexed node, $U_{t}$.
-Specifically, for any nonstationary time-series $i$, we anticipate a connection between $V_{i,t}$ and $U_{t}$ across all time points.
+Specifically, for any nonstationary time series $i$, we anticipate a connection between $V_{i,t}$ and $U_{t}$ across all time points.
 We start from a full graph including the time-indexed node, where all the edges are drawn out.
@@
 This process figures out the edges between nodes $U_{t}$ and $V_{i,t}$.
+The null hypothesis is that the pair are conditionally independent, implying that their conditional joint probability equals the product of their conditional marginals.
+If the obtained p-value falls below the user specified confidence level $\alpha$, signifying dependence, the edge between the pair remains.
 The hypothesis test is:
-$H_{0}\mathrel{\mathop{:}}P(V_{i,t},U_{t}|S)=P(V_{i,t}|S)\,P(U_{t}|S)\,,$
-$H_{1}\mathrel{\mathop{:}}P(V_{i,t},U_{t}|S)\neq P(V_{i,t}|S)\,P(U_{t}|S)\,,$
+$H_{0}:P(V_{i,t},U_{t}|S)=P(V_{i,t}|S)\,P(U_{t}|S)\,,$
+$H_{1}:P(V_{i,t},U_{t}|S)\neq P(V_{i,t}|S)\,P(U_{t}|S)\,,$
 which we test for all subsets $S\subseteq\{V_{k,t-l}\ |\ k=1,...,N;l=0,...,L\}\setminus\{V_{i,t}\}$ where $L$ denotes the maximum considered lag.
 Note that this hypothesis examines the relationship between the time-indexed node and other nodes at the identical time point $t$, maintaining consistency across time $t-l$.
-Afterwards, utilizing the same test, we evaluate the independence of node pairs $V_{i,t-l}$ and $V_{j,t}$ excluding $U_{t}$.
-The null hypothesis is that the pair are conditionally independent, implying that their conditional probability equals the product of their marginals.
-If the obtained p-value falls below the confidence level, signifying dependence, the edge between the pair remains.
-The hypothesis test is:
-$H_{0}\mathrel{\mathop{:}}P(V_{i,t-l},V_{j,t}|S)=P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
-$H_{1}\mathrel{\mathop{:}}P(V_{i,t-l},V_{j,t}|S)\neq P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
+Afterwards, utilizing the same test, we evaluate the independence of node pairs $V_{i,t-l}$ and $V_{j,t}$ excluding $U_{t}$. The hypothesis test is:
+$H_{0}:P(V_{i,t-l},V_{j,t}|S)=P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
+$H_{1}:P(V_{i,t-l},V_{j,t}|S)\neq P(V_{i,t-l}|S)\,P(V_{j,t}|S)\,,$
 where $S\subseteq\{U_{t}\}\cup\{V_{k,t-m}\ |\ k=1,...,N;m=0,...,L\}\setminus\{V_{i,t-l},V_{j,t}\}$.
 It is worth noting that when $l=0$, indicating no lag, we capture the contemporaneous relationships between nodes.
-On the other hand, when $l>0$, we discover the lagged dependencies among the time-series.
+On the other hand, when $l>0$, we discover the lagged dependencies among the time series.
 Under Assumption 4, we maintain the structure consistent over time.
-This means that the discovered dependencies between $V_{i,t-l},V_{j,t}$ remain the same for $V_{i,t-l-m},V_{j,t-m};m=1,...,l-1$.
+This means that the discovered dependencies between $V_{i,t-l},V_{j,t}$ remain the same for $V_{i,t-l-m},V_{j,t-m};m=1,...,L-l$.
 The discovered network is undirected and has the edge minimality condition.
@@
 Stage 3
-Once we have derived the network skeleton, our next step involves orienting the edges based on existing knowledge.
-If there are any nonstationary nodes, meaning there exists an edge between the node and time indexed node, then the orientation is from the time indexed node to the nonstationary node, as time is what causes the node’s corresponding time-series to be nonstationary.
+Once we have derived the network skeleton, our next step involves orienting the edges based on prior knowledge and identifiable structures.
+If there are any nonstationary nodes, meaning there exists an edge between the node and the time-indexed node, then the orientation is from the time-indexed node to the nonstationary node, as time is what causes the node’s corresponding time series to be nonstationary.
 Moreover, all the edges where $l<m$ between pairs $V_{i,t-l}$ and $V_{j,t-m}$, the orientation is from $V_{i,t-m}$ to $V_{j,t-l}$, as logically the effect would happen after the cause.
@@
 In the fourth stage, we rely on the independent changes of the causal modules (Pearl, 2009).
-A dependency measure is defined as an extension of the Hilbert Schmidt Independence Criterion (Gretton et al., 2008), which shows the level of dependency between causal modules.
+A dependency measure is defined as an extension of the Hilbert Schmidt Independence Criterion (Gretton et al., 2007), which shows the level of dependency between causal modules.
 This is done by developing a kernel embedding of nonstationary conditional distributions and using their Gram matrix to create a test statistic (Huang et al., 2017).
 We apply the Meek orientation rule (Meek, 1995) on top of this to orient other edges.
-Note that there may remain edges that are undirected even after the third and fourth stage; in this scenario, no comments can be given on the causal direction of the corresponding pair of nodes.
+Note that this stage is only applicable to the edges between two nonstationary nodes.
+There may remain edges that are undirected even after the third and fourth stages; in this scenario, no comments can be given on the causal direction of the corresponding pair of nodes.
 […]
 Figure 2: CD-NOTS algorithm schematic example
-3.2 Testing the Assumptions
-The main assumptions we claim that are incorrectly assumed in prior causal discovery algorithms is stationarity, no lagged relationships, and linearity.
-Given the fact our algorithm does not make the first two assumptions, we can analyze the graph from our algorithm to see if nonstationary and lagged relationships are found; if found, then the conclusion is that, assuming stationarity (lagged relationships) is the null hypothesis, the statistical tests could not reject the null hypothesis.
-On the other hand, for the linearity assumption, we modify an algorithm used in prior work (Peters et al., 2014) used in order to test post-hoc if the relationships are linear.
-Specifically, the relationships are assumed to be linear in conjunction assuming with additive noise.
+2.2 Testing Assumptions
+In practice, an important step when using algorithms that learn from data is to evaluate (or validate) the results of the algorithm.
+Running interventional studies is one approach; however, this is often infeasible, if not impossible, in fields such as finance.
+Instead, as mentioned in the introduction, evaluation of causal discovery algorithms in the practice rests upon evaluating the realism of the assumptions underlying the algorithm.
+We categorize assumptions into three types:
+  1. 1.
+    Objective Assumptions: These are assumptions which are statistically testable such as stationarity and linearity.
+In Section 2.2.1, we discuss specific ways we can test these two properties.
+Since these assumptions are testable, we argue that any causal discovery algorithms reliant on these assumptions should ensure that the data aligns with them.
+  2. 2.
+    Algorithmic Assumptions (or Algorithmic Hyperparameters): These assumptions refer to decisions in the causal discovery algorithm which can neither be directly evaluated nor justified purely by domain expertise.
+Examples include the significance threshold used in the conditional independence test and the type of conditional independence test.
+While there is some theories and studies that can be used to justify one threshold or test over another (e.g., our simulated experiments in Section 3), we recommend testing the robustness of the causal discovery algorithm’s results to these hyperparameters.
+  3. 3.
+    Subjective Assumptions: These are assumptions which cannot be statistically tested and require domain expertise to evaluate.
+Pseudo-sufficiency, consistency, faithfulness, and randomness (assumptions made by CD-NOTS) fall in this category.
+In our case studies, while we maintain confidence in the validity of these assumptions, the credibility of our algorithm’s results comes from the reader’s belief in these assumptions.
+We discuss the intuition behind these assumptions in Section 2.2.2.
+2.2.1 Objective Assumptions
+One of the main assumptions in prior causal discovery algorithms is stationarity, which can result in serious inaccuracies when dealing with inherently nonstationary data, such as those commonly encountered in fiance and economics.
+Since our algorithm does not rely on this assumption, we can analyze the graph identified by our algorithm to see if nonstationary relationships are found.
+While many recent algorithms do not assume linearity, for comprehensive analysis, we test this assumption post-hoc in our case studies (Section 4) using a modified version of a previously employed algorithm (Peters et al., 2014).
+Typically, prior work assuming linear relationships also assume additive noise.
+Accordingly, when we test for linearity, we also test for additive noise.
 The algorithm used in Peters et al.
-(2014) is to perform a regression and then an independence test on the noise.
-Say the null hypothesis we are testing is that $X$ and $Y$ have a linear relationship given $Z$ (where we allow for $Z$ to have a non-linear relationship with $Y$ and the noise to be heteroscedastic with respect to $Z$), we perform a Gaussian process regression $\hat{f}$, where we use the linear kernel for $X$ and a non-linear kernel (e.g., RBF) for $Z$.
-Having done that, we then perform a CI test on $Y-\hat{f}(X,Y)\perp\!\!\!\perp X|Z$.
+(2014) involves performing a regression followed by an independence test on the noise.
+For instance, if the null hypothesis we are testing is that $X$ and $Y$ have a linear relationship given $Z$ (where we allow for $Z$ to have a non-linear relationship with $Y$ and the noise to be heteroscedastic with respect to $Z$), we perform a Gaussian process regression $\hat{f}$ using a linear kernel for $X$ and a non-linear kernel (e.g., RBF) for $Z$.
+Subsequently, we perform a CI test on $Y-\hat{f}(X,Y)\perp\!\!\!\perp X|Z$.
 Note that our test is more generic than testing if all relationships are linear.
+2.2.2 Subjective Assumptions
+Almost every causal discovery algorithm is built upon subjective assumptions.
+It is crucial to justify these assumptions due to their subjective nature.
+Here, we provide an intuitive explanation of these assumptions in order to allow readers to judge the realism of these assumptions.
+Pseudo-sufficiency
+Any potential confounder described as a smooth function of time can be captured by the time-indexed node; while, if there exists a confounder which is stationary and not changing through time, the algorithm will not capture it.
+Consistency
+The consistency assumption states that the causal links remain unchanged through time.
+We assert any deviations observed in the causal graph over time in Section 4 can be attributed to sampling variations and data availability.
+Faithfulness
+This assumption can be violated in many ways, including xor connections (or gating mechanisms), deterministic functions, or cancelling paths (Marx et al., 2021).
+In our case studies, we do not have situations of deterministic functions which can be directly tested from the data.
+Randomness
+The randomness assumption comes from the CI tests which enables the derivation of the null distribution (statistical validation).
+We note that this assumption does not hold for time series [^3] This is a limitation for all constraint-based causal discovery algorithms that use p-values of CI tests to define a threshold, and this is a gap to be filled in future work.; this implies that the p-value threshold used does not accurately control the false positive rate.
+We argue that testing the robustness of our algorithm to different p-value thresholds mitigates this concern.
 2.3 Importance of Modeling Lagged Relationships
@@
 $X_{0}$ $=0\,.$
+where $W_{t}$ is a Wiener process.
 To discretize, we set $t$ to be integers. Since this is simply an Ornstein-Uhlenbeck (OU) process, we have:
-$P(x_{t}|x_{u})$ $=P(x_{t}|\ \mathcal{N}(x_{u}e^{-\mathinner{\!\left\lvert t-u\right\rvert}},e^{-\mathinner{\!\left\lvert t-u\right\rvert}}))\,,\,u<t\,.$
+[…]
 From this, we can see that $P(X_{t}|X_{0})$ is a time dependent distribution whereas $P(X_{t}|X_{t-1})$ is not.
@@
 […]
-Figure 3: Diagram of example causal graph for which CD-NOD’s time-series algorithm would make a mistake.
+Figure 3: Diagram of example causal graph for which CD-NOD’s time series algorithm would make a mistake.
 3 Evaluation
-Figure 4: F-score evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
-Figure 5: Runtime evaluation for CD-NOTS for different CI tests. Note the y-axis is in log scale.
-In our algorithm, one major design decision is the conditional independence test.
-In order to ascertain which conditional independence test might be most effective, we test our algorithm on different simulated datasets.
-In total, we tested graphs with 3, 4, 5, 6, 8, 10, and 15 nodes, for a total of 350 graphs (50 graphs per number of nodes).
+One major design decision in any constrained-based algorithms, including ours, is the CI test.
+In order to ascertain which test might be most effective, we test our algorithm on different simulated datasets.
+In total, we tested graphs with 3, 4, 5, 6, 8, 10, and 15 nodes, for a total of 350 randomly generated graphs (50 graphs per number of nodes).
 For each graph, we tested having 50, 150, 300, 500, and 1,000 data points.
-For the conditional independence tests, we specifically test using partial correlations (ParCorr, a conditional independence test that assumes linearity), Kernel-based Conditional Independence Test (KCIT), Randomized Conditional Correlation Test (RCoT), and Conditional Mutual Information K-nearest neighbor (CMIknn) Test.
+The relationship between any connected nodes can be either linear or non-linear (quadratic, exponential, or sine).
+The max lag we allowed was five.
+For the conditional independence tests, we specifically test using partial correlations (ParCorr, a conditional independence test that assumes linearity), KCIT, RCoT , and CMIknn.
 For both KCIT and RCoT, there are a few variations in how the p-value can be computed, specifically using the Satterthwaite–Welch (SW) method (Welch, 1938; Satterthwaite, 1946) or Hall–Buckley–Eagleson (HBE) method (Hall, 1983; Buckley and Eagleson, 1988).
@@
 The hyperparameters used for each test can be found in Appendix A.
-In Figure 4, we compare the average F-score for each CI test across a varying number of nodes in the graph and a varying number of data points per graph.
+In Figure 4, we compare the average F-score [^4] F-score is the harmonic mean of precision and recall metrics i.e., F-score$=2\,\frac{precision\times recall}{precision+recall}\,$.
+for each CI test across a varying number of nodes in the graph and a varying number of data points per graph.
 As expected, the F-score improves as the number of data points increase.
 Possibly most interesting is that ParCorr performs the best for low data regimes.
-Further, we can see, as found by previous work (Bodenham and Adams, 2015), that the SW variant of KCIT tends to be weaker than HBE (except when there are a large number of nodes and only 50 data points).
+Further, we can see, as found by previous work (Bodenham and Adams, 2015), that the SW variant of KCIT tends to be weaker than HBE (except when there is a large number of nodes and only 50 data points).
 However, there seems to be little difference for RCoT.
-So, relying on results in Bodenham and Adams (2015), we recommend the HBE variant.
-Finally, while CMIknn is quite competitive, we do not recommend its usage due to its lengthy runtime (Figure 5).
+Hence, relying on results in Bodenham and Adams (2015), we recommend the HBE variant.
+Figure 4: F-score evaluation for CD-NOTS with different CI tests, tested out on different simulated datasets.
+From our simulation studies, we recommend using RCoT for large datasets since it is much faster than KCIT on a CPU.
+However, we do note that the difference in speed is much less when using a GPU implementation of KCIT and RCoT; thus, if a GPU is available, we recommend KCIT since it does not have a randomized component.
+Finally, while CMIknn is quite competitive, we do not recommend its usage on large problems due to its lengthy runtime.
+Another important metric to consider while choosing a CI test is the computational cost of it.
+Figure 5 shows the average runtime of different tests for different sample sizes.
+In general, methods like CMIKnn are several orders of magnitude slower than simple linear tests like ParCorr which limits their applicability in very large samples.
+In terms of computations, KCIT is positioned in the middle and its cost grows in $\mathcal{O}(n^{3})$ with sample size.
+RCoT has time complexity of $\mathcal{O}(m^{2}n)$ where $m$ is the number of Fourier terms (typically in order of 10-100) set by the user which gives it a considerable computation saving when compared to KCIT.
+In general, as expected the number of conditions that needs to be tested grows rapidly with the number of nodes as shown in Figure 6.
+Note that the number of tests needed for a given network size will not be exactly the same for different CI tests due to fact that any edge removal in early states of skeleton discovery (which depends on the p-value generated by the given test) will result in a substantial decrease in condition sets that need to be tested and hence the total number of tests.
+Figure 5: Runtime evaluation for CD-NOTS for different CI tests. Note the y-axis is in log scale.
+Figure 6: Number of tests run for CD-NOTS for different CI tests. Note the y-axis is in log scale.
+Another important parameter of the algorithm is the significance level $\alpha$.
+We ran a series of experiments with $\alpha$ levels [0.01, 0.05, 0.1] to study the sensitivity of the algorithm to the choice of $\alpha$.
+The results for F-score have been shown in Figure 7.
+Please see Appendix D.2 for other metrics.
+As expected, a lower $\alpha$ results in higher precision across all data ranges and CI tests.
+The only exception to this is RCoT in low data regimes, but this is expected because RCoT, by design, assumes there are large number of data points and its results are not reliable for small datasets.
+In terms of F-score, larger $\alpha$ values perform better in low data regimes; however, in large data regimes, small alphas perform better.
+(a) ParCorr
+(b) CMIKnn
+(c) KCIT
+(d) RCoT
+Figure 7: F-score evaluation, comparing different confidence levels.
 Comparing against the best tests (KCIT HBE, RCoT HBE, ParCorr, CMIknn), we consider PCMCI (Runge et al., 2019) as a benchmark since it is a comparable PC-based algorithm and is widely used in the causal discovery literature.
@@
 4 Case Studies
-We apply CD-NOTS on three different datasets: 1) Fama-French factors and Apple’s returns, 2) unemployment, CPI, and PPI of different countries, and 3) Price-to-Book ratio and stock returns of financial companies in the S&P 500.
 Most of the work done in factor investing and economic indices’ analysis rely on the correlation between the variables and do not consider the fundamental question of “why” (López de Prado, 2023).
-In each case study, we explain the data and properties we see in the data or has been found in prior work, discuss the implications of the causal networks found by CD-NOTS, and, finally, show that the assumptions of stationarity, linearity, and no lagged relationships would be incorrect to assume and thus, any algorithm making said assumptions would be inaccurate and invalid to use on these datasets.
-For all the graphs in the section, we color by the maximum p-value for that edge across all tests run during the causal skeleton step (Stage 2).
+We apply CD-NOTS to three financial datasets: 1) Fama-French factors (Fama and French, 1993) and Apple’s returns, 2) unemployment, CPI, and PPI of different countries, and 3) Price-to-Book ratio and stock returns of financial companies in the S&P 500.
+In each case study, we explain the data and properties we see in the data or that has been found in prior work, discuss the implications of the causal networks found by CD-NOTS and, finally, show that the assumptions of stationarity, linearity, and no lagged relationships would be incorrect to assume and thus, any algorithm making said assumptions would be inaccurate to use on these datasets.
+For all the graphs in this section, we color by the maximum p-value for that edge across all tests ran during the causal skeleton step (Stage 2).
 While we recognize that p-values are not a metric to measure the strength of the relationship, we can think of the vicinity of the p-value to 0.05 as a measure of the sensitivity of the graph to that hyperparameter (i.e., the threshold).
+While in our experiments in Section 3 we found that using a p-value of 0.01 to be better for larger datasets, we run our experiments with a p-value of 0.05, both to be consistent with prior work (Strobl et al., 2017) as well as to show the edges that would be removed by changing from 0.05 to 0.01.
 4.1 Fama-French Factors
-Figure 7: Time-series of Fama-French factors and Apple’s returns.
-Refer to caption Refer to caption Refer to caption Refer to caption (a) Jan 2000 - Dec 2007 Refer to caption (b) Jan 2008 - Dec 2015 Refer to caption (c) Jan 2016 - Dec 2022
-Figure 8: Results of CD-NOTS KCIT run on three non-overlapping time periods.
-Refer to caption Refer to caption
-Figure 9: Results of CD-NOTS KCIT run from beginning of 2000 to end of 2022.
 Data
-For our first case study, we analyzed the relationship between the Fama-French factors (Fama and French, 1993) and Apple’s returns from the beginning of 2000 to the end of 2022.
-In the data in Figure 7, we show that there is some correlation between the variables (e.g., the spike in volatility around 2008 can be seen in all six subplots) and, accordingly, all variables show volatility clustering.
+For our first case study, we analyze the relationship between the Fama-French factors and daily Apple’s returns from the beginning of 2000 to the end of 2022.
+In the data in Figure 9, we show that there is some correlation between the variables (e.g., the spike in volatility around 2008 and during COVID Pandemic can be seen in all six subplots) and, accordingly, all variables show volatility clustering.
+Figure 9: Time series of Fama-French factors and Apple’s returns.
 Results and Discussion
-In Figure 8, we ran CD-NOTS using KCIT (RCoT results are in Appendix B) on three non-overlapping periods of six to seven years (approximately 2000 observations each).
-We note in all three periods that there is no connection from $T$ to Apple’s returns (AAPL_RET), implying that the nonstationarity observed in Apple’s returns can be explained away by the Fama-French factors.
-Further, the lagged relationships tends to be sparse, with many of the found lagged relationships having a high p-value (close to 0.05).
-However, we see that the relationship between the factors seems to change over time, as well as the relationship between the factors and Apple’s returns.
-In Figure 9, we run CD-NOTS with KCIT on the whole time range, and we can see now that that there is a connection from $T$ to Apple’s returns (AAPL_RET).
-This is to be expected for two reasons: 1) over the course of the last 22 years, Apple’s business model has changed, leading to exposures to different factors over time, and 2) the changing relationship between the factors and Apple’s returns, as seen in Figure 8.
+In Figure 10 and Figure 11, we run CD-NOTS with KCIT on the whole time range with and without the $T$ node, respectively, or, in other words, test the impact of modeling nonstationarity.
+While the number of edges in the contemporaneous graph is about the same (one edge difference, specifically between SMB and CMA), there are fewer lagged relationships in Figure 10.
+This implies the dependence we see through time is better explained as a nonstationary relationship.
+It should be noted that there is a connection from $T$ to Apple’s returns (AAPL_RET).
+This is to be expected for two reasons: 1) over the course of the last 22 years, Apple’s business model has changed, leading to exposures to different factors over time, and 2) the changing relationship between the factors and Apple’s returns.
+Figure 10: Results of CD-NOTS KCIT run from beginning of 2000 to end of 2022. Left Contemporaneous graph Right Full time series graph
+Figure 11: Results of CD-NOTS KCIT run from beginning of 2000 to end of 2022 without the time-indexed node. Left Contemporaneous graph Right Full time series graph
 Evaluating Assumptions
-One of the assumptions of CD-NOTS is causal consistency that the causal graph does not change over time.
-We assert that the change in dependencies between the factors and Apple’s returns that we see in Figure 8 are not a result of changes in the causal directions but in the strength of the causal relationships as captured by the nonstationarity.
-Further, we note that the fact that there is a relationship between $T$ and all the Fama-French factors (nonstationarity) implies that any causal discovery algorithm assuming stationarity would give invalid results.
+We note that the fact that there is a relationship between $T$ and all the Fama-French factors (nonstationarity) implies that any causal discovery algorithm assuming stationarity would give invalid results.
 Similarly, we find statistically significant lagged relationships.
@@
 4.2 Economic Data
-Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
-Figure 10: Results of CD-NOTS ParCorr run on six countries’ economic variables.
-Figure 11: Results of CD-NOTS KCIT run on the economic variables of multiple countries.
 Data
-For our second case study, we analyzed the relationship between the month-over-month percentage change in CPI, the month-over-month percentage change in PPI, and unemployment for the U.S., Japan, Canada, India, Italy, the U.K., and France.
+For our second case study, we analyze the relationship between the month-over-month percentage change in Consumer Price Index (CPI), the month-over-month percentage change in Producer Price Index (PPI), and unemployment for the U.S., Japan, Canada, India, Italy, the U.K., and France.
 We use monthly data from the beginning of 2000 to the end of 2023.
+(a) U.S.
+(b) Canada
+(c) Japan
+(d) France
+(e) United Kingdom
+(f) Italy
+Figure 12: Results of CD-NOTS ParCorr ran on six countries’ economic variables.
 Results and Discussion
@@
 Similar to the previous case study in which the graph over all time periods is akin to a superset of the edges of individual time periods, we see a similar behavior here.
+One difference we see is a new link from $T$ to the percentage change in PPI which is not in any country-specific causal graphs; we interpret this as finding a new relationship through increasing the statistical power by pooling the data.
 Note that Country does not point to any other variable suggesting that the behavior across countries is independent of the country; this observation could be used to justify training a single econometric model across countries.
+Figure 13: Results of CD-NOTS KCIT ran on the economic variables of multiple countries.
 Evaluation Assumptions
@@
 We note that this is inconsistent with our usage of ParCorr in Figure 12 since ParCorr is a linear-based CI test.
-In Appendix C, we show the results of using KCIT and RCoT where we find qualitatively that the graphs are similar.
+In Appendix C, we show the results of using RCoT where we find qualitatively that the graphs are similar.
 4.3 Company Financials and Returns
-Refer to caption (a) Contemporaneous Effects Refer to caption (b) Lagged Effects Refer to caption (c) Contemporaneous Effects Refer to caption (d) Lagged Effects Refer to caption (e) Comparison of Returns and N.
-LPTB Refer to caption (f) Cumulative Returns of Strategies
+(a) Contemporaneous Effects
+(b) Lagged Effects
+(c) Contemporaneous Effects
+(d) Lagged Effects
 Figure 14: Results of CD-NOTS KCIT run on financials data for log Price-to-Book ratio (LPTB) and stock returns (Ret).
@@
 LPTB) refers to normalizing the LPTB by the mean and standard deviation of the LPTB of financial companies in that quarter.
-For Figure 11(f), our strategy is long-ing financial companies that have a normalized LPTB greater than the threshold which we compare against the market-cap weighted portfolio.
 Data
-For our third case study, we analyzed the relationship between the Price-to-Book ratio and returns of financial companies in the S&P 500.
+For our third case study, we analyze the relationship between the Price-to-Book ratio and returns of financial companies in the S&P 500.
 We use data from the beginning of 2010 to the end of 2023.
@@
 To ensure no look-ahead bias when defining the Price-to-Book ratio, for any given date (i.e., end of quarter), we use the Price-to-Book ratio available that day.
-We join the next quarter returns with this company financials information.
+We join the next quarter returns with this company’s financial information.
 For our experiments, we use the log Price-to-Book ratio (LPTB) as well as the normalized LPTB, which we define as the LPTB normalized by the mean and standard deviation of the LPTB of financial companies in that quarter.
@@
 LPTB).
-Further, we see that the previous quarter’s normalized LPTB point to the next quarter’s returns with a reasonable p-value suggesting that its inclusion in the graph is not purely a function of the threshold we chose.
-In Figure 11(e), we compare the next quarter returns and the normalized LPTB and see that the there is a small trend in the mean (the black circles).
+Further, we see that the previous quarter’s normalized LPTB points to the next quarter’s returns with a reasonable p-value suggesting that its inclusion in the graph is not purely a function of the threshold we chose.
+In Figure 15(a), we compare the next quarter returns and the normalized LPTB and see that there is a small trend in the mean (the black circles).
 To this end, we create a simple trading strategy where we invest in the subset of the financial securities whose LPTB is greater than some threshold.
-In Figure 11(f), we see this simple approach outperforms a market-cap weighted strategy; prior work (Monge et al., 2023) has found that growth stocks have been outperforming value stocks.
+In Figure 15(b), we see this simple approach outperforms a market-cap weighted strategy; prior work (Monge et al., 2023) has found that growth stocks have been outperforming value stocks.
 Importantly, we note that the goal of this is to not introduce a new signal but to show how causal discovery can be a tool in searching for statistically significant relations which, if the assumptions are true for the data, can be interpreted as causal.
+(a) Comparison of Returns and N. LPTB
+(b) Cumulative Returns of Strategies
+Figure 15: Testing out predictive power of Normalized LPTB for returns of financial companies.
+For Figure 15(b), our strategy is long-ing financial companies that have a normalized LPTB greater than the threshold which we compare against the market-cap weighted portfolio.
 Evaluating Assumptions
 Similar to previous case studies, we see statistically significant lagged relationships and statistically significant nonstationarity, implying we would not be able to trust the results of an algorithm that cannot handle these types of properties.
-For our linearity test, we test the linearity of the normalized LPTB and returns conditional on all other parent nodes of returns and find the p-value be to 0.0 (i.e., we can reject the null hypothesis that the relationship is linear).
+For our linearity test, we test the linearity of the normalized LPTB and returns conditional on all other parent nodes of returns and find the p-value to be 0.0 (i.e., we can reject the null hypothesis that the relationship is linear).
 5 Conclusion and Future Work
-In this paper, we developed a novel algorithm, Constraint-based Causal Discovery from Heterogeneous Time-Series (CD-NOTS), which is nonparametric (meaning it can capture both linear and non-linear causal relations), can handle nonstationarity, and is able to detect both lagged and contemporaneous causalities.
+In this paper, we developed a novel algorithm, Constraint-based Causal Discovery from Nonstationary Time Series (CD-NOTS), which is nonparametric, can handle nonstationarity, and is able to detect both lagged and contemporaneous causalities.
 We showed using simulated data that, for effective use of CD-NOTS, one should use ParCorr for low data regimes (less than 200 data points) and KCIT or RCoT for higher data regimes.
-We applied CD-NOTS to Fama-French factors and Apple’s returns, where we found that for small time ranges (approximately seven years), Apple’s returns are stationary conditioning on the factors.
-This justifies the usage of Fama-French factors as a method for returns attribution.
-However, over longer time ranges (approximately two decades), we find Apple’s returns are nonstationary even after accounting for the Fama-French factors, implying that there is a need to continually update the factor exposures of Apple through time.
+We applied CD-NOTS to approximately two decades of Fama-French factors and Apple’s returns and found that Apple’s returns are nonstationary even after accounting for the Fama-French factors, implying that there is a need to continually update the factor exposures of Apple through time.
 Further, we found the causal relationship of economic factors (specifically unemployment, CPI, and PPI) tends to be similar across countries, especially for those geographically and economically similar.
@@
 Finally, we apply CD-NOTS to company financials and stock returns and find a causal relationship between the Price-to-Book ratio and future returns of financial companies, justifying the common usage of Price-to-Book ratio for investing.
-Besides our contribution in developing an algorithm that can handle nonstationarity, non-linear relationships, lagged and contemporaneous causalities, we showed the importance of these assumptions in finance.
-Specifically, many of the assumptions made in prior work (notably linearity, no lagged relationships, and stationarity) often do not hold in finance and thus should be used with caution.
-Hence, through weakening the assumptions required for our causal discovery algorithm to be used, we can trust more the causal relationships found by CD-NOTS and continue to explore the causal relationships in other financial datasets.
+Through our case studies, we show the importance of testing the assumptions of causal discovery algorithms.
+Specifically, we find in all of our case studies nonstationarity, non-linear relationships, and lagged and contemporaneous causal effects.
+Prior work often assumes some of these properties are not present in the data and thus would lead to spurious relationships and incorrect conclusions.
+Hence, our contribution improves over prior work by handling all of these scenarios.
 Appendix A Conditional Independence Tests Hyperparameters
-For all conditional independence tests, we use a threshold of 0.05 for the p-value.
+For all conditional independence tests in our case studies, we use a threshold of 0.05 for the p-value.
 KCIT
@@
 We repeat the plots from our first case study (Section 4.1) using RCoT as opposed to KCIT.
-One empirical finding was that, due to the randomness of RCoT, many times there could be an inconsistency between the discovered graph and the CI tests, i.e., the graph shows conditional independences that are not found by the CI tests.
-We leave it to future work to find how to account for these forms on inconsistencies.
-Refer to caption Refer to caption Refer to caption Refer to caption (a) Jan 2000 - Dec 2007 Refer to caption (b) Jan 2008 - Dec 2015 Refer to caption (c) Jan 2016 - Dec 2022
-Figure 13: Results of CD-NOTS RCoT run on three non-overlapping time periods.
-Refer to caption Refer to caption
+One empirical finding is that, due to the randomness of RCoT, many times there could be an inconsistency between the discovered graph and the CI tests, i.e., the graph shows conditional independences that are not found by the CI tests.
+We leave it to future work to find how to account for these forms of inconsistencies.
 Figure 16: Results of CD-NOTS RCoT run from beginning of 2000 to end of 2022.
 Appendix C Economic Data Case Study with KCIT and RCoT
-Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
-Figure 15: Results of CD-NOTS KCIT run on six countries’ economic variables.
-Refer to caption (a) U.S. Refer to caption (b) Canada Refer to caption (c) Japan Refer to caption (d) France Refer to caption (e) United Kingdom Refer to caption (f) Italy
-Figure 16: Results of CD-NOTS RCoT run on six countries’ economic variables.
-Figure 17: Results of CD-NOTS RCoT run on the economic variables of multiple countries.
+Our synthetic data experiments indicate the effectiveness of ParCorr when data is limited, as observed in our country-level experiments.
+However, in this section, we present results obtained using KCIT and RCoT.
+Notably, these results appear to lack robustness to specific test used, emphasizing the importance of analyzing countries collectively.
+Upon comparing KCIT with RCoT, we note several similarities in edges, barring a few exceptions.
+For instance, RCoT suggests that Unemployment and PPI are stationary, whereas CPI lacks a one-lag relationship with itself.
+Additionally, it indicates an additional edge from Unemployment in the previous month to the current month’s CPI.
+(a) U.S.
+(b) Canada
+(c) Japan
+(d) France
+(e) United Kingdom
+(f) Italy
+Figure 17: Results of CD-NOTS KCIT ran on six countries’ economic variables.
+(a) U.S.
+(b) Canada
+(c) Japan
+(d) France
+(e) United Kingdom
+(f) Italy
+Figure 18: Results of CD-NOTS RCoT ran on six countries’ economic variables.
 Appendix D Detailed Simulation Results
+Further findings on simulated data includes an evaluation of precision, recall555precision$=\frac{tp}{tp+fp}\,$, recall$=\frac{tp}{tp+fn}$ where $tp=$true positive, $fp=$false positive and $fn=$ false negative.
+and Structural Hamming Distance (SHD).
+Note that while precision, recall, and F-score aim for maximization, the objective for SHD is minimization.
+Figure 19: Results of CD-NOTS RCoT ran on the economic variables of multiple countries.
 D.1 Different CI Tests
-Figure 18: F-score evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
-Figure 19: Precision evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
-Figure 20: Recall evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
-Figure 21: Structural Hamming Distance evaluation for CD-NOTS with different CI tests, tested out on many different simulated datasets.
-D.2 Comparison against PCMCI
-Figure 22: F-score evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
-Figure 23: Precision evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
-Figure 24: Recall evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
-Figure 25: Structural Hamming Distance evaluation, comparing CD-NOTS and PCMCI, tested out on many different simulated datasets.
+At 300 data points, KCIT-SW demonstrates the highest recall but substantially lower precision, resulting in the poorest F-score within the plot.
+Conversely, in the context of 1000 data points, recall remains consistent across all methods, with discrepancies in F-score attributed to variations in precision.
+Notably, with 50 data points, the precision curve exhibits an unexpected upward trend, contrary to the anticipated decrease in precision (and recall) as the node count increases.
+Interestingly, despite these nuances, SHD yields a performance ranking similar to that of F-score across the evaluated scenarios.
+Figure 20: Precision evaluation for CD-NOTS with different CI tests, tested out on different simulated datasets.
+Figure 21: Recall evaluation for CD-NOTS with different CI tests, tested out on different simulated datasets.
+Figure 22: SHD evaluation for CD-NOTS with different CI tests, tested out on different simulated datasets.
+D.2 Effect of $\alpha$
+Note that the significance level $\alpha$ controls the false positive rate.
+i.e., when employing a conditional independence test, if the data is truly independent, the test will fail $\alpha$ percent of the time, yet it does not explicitly address false negatives.
+This phenomenon becomes evident when observing precision plots, where the optimal precision corresponds to the lowest $\alpha$, indicating the lowest false positive rate.
+Conversely, recall operates inversely, as a decrease in false positives generally coincides with an increase in false negatives.
+It is this balance that the F-score metric achieves, although it is worth noting that F-score is not the sole method to combine the two metrics together.
+Regarding SHD, it presents another perspective on balancing false positives and false negatives.
+Unlike F-score, SHD typically advocates for an alpha of 0.01, except for specific scenarios, such as when dealing with 50 and 300 data points in RCoT and CMIknn for 300 data points with larger graphs.
+(a) ParCorr
+(b) CMIKnn
+(c) KCIT
+(d) RCoT
+Figure 23: Effect of $\alpha$ on Precision
+(a) ParCorr
+(b) CMIKnn
+(c) KCIT
+(d) RCoT
+Figure 24: Effect of $\alpha$ on Recall
+(a) ParCorr
+(b) CMIKnn
+(c) KCIT
+(d) RCoT
+Figure 25: Effect of $\alpha$ on SHD
+D.3 Comparison against PCMCI
+CD-NOTS consistently outperforms PCMCI in terms of SHD, similar to its dominance in the F-score metric.
+While PCMCI exhibits better recall than CD-NOTS, it does so at the cost of significantly lower precision.
+This trade-off explains why CD-NOTS achieves better performance in both F-score and SHD metrics.
+Figure 26: Precision evaluation, comparing CD-NOTS and PCMCI, tested out on different simulated datasets.
+Figure 27: Recall evaluation, comparing CD-NOTS and PCMCI, tested out on different simulated datasets.
+Figure 28: SHD evaluation, comparing CD-NOTS and PCMCI, tested out on different simulated datasets.